# Tracr-Injection: Distilling Algorithms into Pre-trained Language Models

### v1

@@ Abstract
Motivated by the surge of large language models, there has been a push to formally characterize the symbolic abilities intrinsic to the transformer architecture. A programming language, called RASP, has been proposed, which can be directly compiled into transformer weights to implement these algorithms. However, the tasks that can be implemented in RASP are often uncommon to learn from natural unsupervised data, showing a mismatch between theoretical capabilities of the transformer architecture, and the practical learnability of these capabilities from unsupervised data. We propose tracr-injection, a method that allows us to distill algorithms written in RASP directly into a pre-trained language model. We showcase our method by injecting 3 different algorithms into a language model. We show how our method creates an interpretable subspace within the model’s residual stream, which can be decoded into the variables present in the code of the RASP algorithm. Additionally, we found that the proposed method can improve out of distribution performance compared to our baseline, indicating that indeed a more symbolic mechanism is taking place in the inner workings of the model. We release the code used to run our experiments [^1] https://github.com/tvergara/tracr-injection.
@@ 1 Introduction
Figure 1: An overview of proposed method: we distill the residual stream of a compiled transformer using tracr, into a pre-train model. We enforce the residual stream to linearly encode the compiled model’s representations.
@@ 3 Methods
where roll shifts activations along the batch dimension (using detach to prevent cross-batch gradients). This encourages the model to utilize the injected symbolic representation subspace, without altering too much the activation distribution.
@@ 5 Experiments
We additionally test the final models on SST-2 Socher et al. (2013) and MRPC Dolan and Brockett (2005) from GLUE Wang (2018), which are binary classification tasks. We select the chosen option as the one with the highest logits. The prompts used are shown in Appendix F. Also, we add an evaluation of the model on LAMBADA Paperno et al. (2016), which is a next word prediction dataset. Finally, we test the perplexity on an unsupervised text dataset that is out-of-distribution with respect to the data used during finetuning, specifically the Tiny-Shakespeare corpus Blog (2015). All of these evaluations are conducted to assess the extent to which the model is able to avoid catastrophic forgetting Kirkpatrick et al. (2017) of the skills it learned during pre-training.
All of our experiments share the same hyperparameters, which are shown in the Appendix B. We run each experiment with the same seed for reproducibility of our results.
@@ 6 Results
Table 2: Performance in out of distribution scenarios comparing Tracr-Injection (TI) and our finetuning + KL baseline (FT). A full description of each setting is in Appendix D.
@@ 7.1 Symbolic Representations Learned
Figure 2: An example of how we can interpret the residual stream of the final model given the final linear layer used during training. Our decoder is given directly from our compiled model, were we know how each dimension in the residual stream maps to a particular value of a variable in the RASP program. An example of this process is found in Appendix E.
@@
We design several out-of-distribution evaluations for each algorithm. Some of these tests are tasks that the compiled RASP model would solve perfectly, while others fall outside the scope of the compiled model due to the nature of the implemented algorithm. We expected tracr-injection to be able to generalize to some extent in out-of-distribution scenarios where the compiled transformer model supported the scenario, while not being able to generalize in other cases.
We show our results in Table 2. We find that our method can enhance the out-of-distribution generalization in some settings. In the case of shuffle-dyck, the accuracy in all out of distribution settings practically did not suffer, while the baseline considerably achieved much lower scores. In other settings, the difference between our proposed method and the baseline is not as strong, perhaps suggesting that the overall out-of-distribution ability is linked to the specific algorithm which is being distilled. We are optimistic that future work more focused on distilling the causal structure from the compiled network Geiger et al. (2022) can have the potential to overcome these limitations and make out-of-distribution capabilities generalize independent of the algorithm at hand.
@@ 8 Future Work
Our method provides the first step towards truly injecting a ground truth mechanism into a pre-trained language model. However, the lack of consistent out-of-distribution generalization highlights that more work needs to be done in order to truly inject a mechanism that generalizes. One avenue of exploration is understanding why certain injected algorithms seem to generalize consistently better than other algorithms (as the shuffle-dyck algorithm). There might be connections between these results and RASP-L Zhou et al. (2023). As RASP-L is a "learnable" subset of RASP instructions, it might be that some algorithms were inherently “harder" to truly learn than others. Additionally, scaling our method to more challenging tasks is a fruitful avenue forward, as all of our tasks could be easily learned by both the baseline and our method. Scaling to these more complex tasks comes with its own complications (as the implementation of ever more complex RASP programs), but it can highlight the benefits of using Tracr-Injection in comparison to finetuning.
@@ 9 Limitations
While our methods do provide a way of distilling the RASP variables into the residual stream of a transformer, the out of distribution tests suggests the model is not representing these same variables when tested in examples which are radically outside of the distribution seen in training. This is a big limitation, as a fully injected algorithm should work in all distributions of inputs, and not just in a vicinity. Additionally, the use of 3 simple symbolic tasks to showcase our method is another limitation. The injection of more complex RASP programs can potentially highlight the performance benefits of our method in constrast to finetuning. The final limitation we mention is the focus exclusively on GPT-2 to evaluate the performance of our work. Compute is an important restriction for which we have made this decision, and we encourage further work in scaling up our methods into more (and larger) language models.
@@ Appendix C Details on each implemented task
Appendix D Out of distribution examples
Bellow we describe the examples we used in each of our evaluations on out of distribution settings. We chose 3 settings for each task, with one of these settings being a "tracr-supported" setting, which means that the original tracr-compiled model from the RASP code would solve this setting perfectly.
@@ Appendix D Out of distribution examples
    • X-only data: we use 40 examples with only "x" values, so for example "x x x x x", from 1 to 40 "x"s (as it was train to only see sequences up to 40). It has to count every single character. These examples were not seen in training. The compiled model can solve this scenario.
@@
    • Cascading Overflow: We test it on data which the model has to overflow digits all the way from the right to the left, for example 299 + 1 (the model has to push the overflow digit all the way to the beggining). It is a main difficulty in implementing a sum algorithm in RASP, so we expected our method to work better. The compiled model can solve this scenario.
@@
    • Additional parentheses family: We add the parentheses family "<>" which is unsee in the training data. Our compiled model does not solve this scenario, as only 3 parentheses families where specified in the compilation ("()", "{}", "[]").

### v2

@@ Abstract
Motivated by the surge of large language models, there has been a push to formally characterize the symbolic abilities intrinsic to the transformer architecture. A programming language, called RASP, has been proposed, which can be directly compiled into transformer weights to implement these algorithms. However, the tasks that can be implemented in RASP are often uncommon to learn from natural unsupervised data, showing a mismatch between theoretical capabilities of the transformer architecture, and the practical learnability of these capabilities from unsupervised data. We propose tracr-injection, a method that allows us to distill algorithms written in RASP directly into a pre-trained language model. We showcase our method by injecting 3 different algorithms into a language model. We show how our method creates an interpretable subspace within the model’s residual stream, which can be decoded into the variables present in the code of the RASP algorithm. Additionally, we found that the proposed method can improve out-of-distribution performance compared to our baseline, indicating that indeed a more symbolic mechanism is taking place in the inner workings of the model. We release the code used to run our experiments [^1] https://github.com/tvergara/tracr-injection.
@@ 1 Introduction
Figure 1: An overview of proposed method: we distill the residual stream of a compiled transformer using tracr, into a pre-trained model. We enforce the residual stream to linearly encode the compiled model’s representations.
@@ 3 Methods
where roll shifts activations along the batch dimension (using detach to prevent cross-batch gradients). This encourages the model to utilize the injected symbolic representation subspace, without altering the activation distribution too much.
@@ 5 Experiments
To measure whether our method introduces catastrophic forgetting Kirkpatrick et al. (2017), we evaluate the final models on tasks that probe both classification and language-modeling abilities:
Classification retention
We run inference on the binary GLUE Wang (2018) tasks SST-2 Socher et al. (2013) and MRPC Dolan and Brockett (2005), selecting the class with the highest logit.
Next-word prediction
We measure accuracy on LAMBADA Paperno et al. (2016), which tests long-range contextual understanding.
Language-model perplexity
We compute perplexity on Tiny-Shakespeare Blog (2015), an unsupervised corpus disjoint from the finetuning data.
Only the perplexity measured on the o.o.d. Tiny-Shakespeare corpus appears in the main text (Table 1); the remaining results are deferred to Appendix (Table 3). All of our experiments share the same hyperparameters, which are shown in the Appendix B. We run each experiment with the same seed for reproducibility of our results.
@@ 6 Results
Table 2: Performance in out-of-distribution scenarios comparing Tracr-Injection (TI) and our finetuning + KL baseline (FT). A full description of each setting is in Appendix D.
@@ 7.1 Symbolic Representations Learned
Figure 2: An example of how we can interpret the residual stream of the final model given the final linear layer used during training. Our decoder is given directly from our compiled model, where we know how each dimension in the residual stream maps to a particular value of a variable in the RASP program. An example of this process is found in Appendix E.
@@
We design several out-of-distribution (o.o.d.) evaluations for each algorithm. Some of these tests are tasks that the compiled RASP model would solve perfectly, while others fall outside the scope of the compiled model due to the nature of the implemented algorithm. We expect Tracr-injection to be able to generalize to some extent in o.o.d. scenarios where the compiled transformer model supported the scenario, while not being able to generalize in other cases.
We show our results in Table 2. We find that our method can enhance the o.o.d. generalization in some settings. In the case of shuffle-dyck, the accuracy in all out-of-distribution settings practically did not suffer, while the baseline considerably achieved much lower scores. In other settings, the difference between our proposed method and the baseline is not as strong, perhaps suggesting that the overall o.o.d. ability is linked to the specific algorithm which is being distilled. We are optimistic that future work more focused on distilling the causal structure from the compiled network Geiger et al. (2022) can have the potential to overcome these limitations and make o.o.d. capabilities generalize independent of the algorithm at hand.
@@ 8 Future Work
Our method provides the first step towards truly injecting a ground truth mechanism into a pre-trained language model. However, the lack of consistent o.o.d. generalization highlights that more work needs to be done in order to truly inject a mechanism that generalizes. One avenue of exploration is understanding why certain injected algorithms seem to generalize consistently better than other algorithms (as the shuffle-dyck algorithm). There might be connections between these results and RASP-L Zhou et al. (2023). As RASP-L is a "learnable" subset of RASP instructions, it might be that some algorithms were inherently “harder" to truly learn than others. Additionally, scaling our method to more challenging tasks is a fruitful avenue forward, as all of our tasks could be easily learned by both the baseline and our method. Scaling to these more complex tasks comes with its own complications (as the implementation of ever more complex RASP programs), but it can highlight the benefits of using Tracr-Injection in comparison to finetuning.
@@ 9 Limitations
While our methods do provide a way of distilling the RASP variables into the residual stream of a transformer, the out-of-distribution tests suggests the model is not representing these same variables when tested in examples which are radically outside of the distribution seen in training. This is a big limitation, as a fully injected algorithm should work in all distributions of inputs, and not just in a vicinity. Additionally, the use of 3 simple symbolic tasks to showcase our method is another limitation. The injection of more complex RASP programs can potentially highlight the performance benefits of our method in constrast to finetuning. The final limitation we mention is the focus exclusively on GPT-2 to evaluate the performance of our work. Compute is an important restriction for which we have made this decision, and we encourage further work in scaling up our methods into more (and larger) language models.
10 Acknowledgements
This project was funded by the National Center for Artificial Intelligence CENIA FB210017, Basal ANID.
@@ Appendix C Details on each implemented task
Appendix D Out-of-distribution examples
Below we describe the examples we used in each of our evaluations on out-of-distribution settings. We chose 3 settings for each task, with one of these settings being a "tracr-supported" setting, which means that the original tracr-compiled model from the RASP code would solve this setting perfectly.
@@ Appendix D Out-of-distribution examples
    • X-only data: we use 40 examples with only "x" values, so for example "x x x x x", from 1 to 40 "x"s (as it was trained to only see sequences up to 40). It has to count every single character. These examples were not seen in training. The compiled model can solve this scenario.
@@
    • Cascading Overflow: We test it on data which the model has to overflow digits all the way from the right to the left, for example 299 + 1 (the model has to push the overflow digit all the way to the beginning). It is a main difficulty in implementing a sum algorithm in RASP, so we expected our method to work better. The compiled model can solve this scenario.
@@
    • Additional parentheses family: We add the parentheses family "<>" which is unseen in the training data. Our compiled model does not solve this scenario, as only 3 parentheses families where specified in the compilation ("()", "{}", "[]").